OpenStudy (anonymous):

If the system y=x^2+2x+3 and y=kx-1 has exactly one solution, then what does k equal?

OpenStudy (anonymous):

\[y = x^{2} + 2x + 3\] \[y = kx - 1\] Since they have one solution, set the equal to each other.. \[x^{2} + 2x + 3 = kx - 1\]

OpenStudy (anonymous):

Will K have a value in this instance or will it be: k = x+2+4/x ?

OpenStudy (anonymous):

\[x^{2} - (k - 2)x + 4 = 0\] \[x^{2} - 4x + 4 = 0\] The second one gives 1 solution. Do you see what K should be?

OpenStudy (anonymous):

I understand how you got the first equation but not sure how you got to the second one.

OpenStudy (anonymous):

That one has no correlation with the first equation, it is there because it has one solution and has the same format as the first one (x^2 ... -4)

OpenStudy (anonymous):

Ah I think I understand, just bear with me, K should be 2?

OpenStudy (anonymous):

Not quite Type out what is the ONLY difference between the 2 equations and post it here

OpenStudy (anonymous):

Okay,(k-2)x and (4x)

OpenStudy (anonymous):

Now, take out the x and what do you have?

OpenStudy (anonymous):

k-2 and 4

OpenStudy (anonymous):

Right, now set these 2 equal and see what you get.

OpenStudy (anonymous):

K=6 so when you multiply out (6-2)x it equals 4x so the equations will be the same, I got it! I think..

OpenStudy (anonymous):

Just put your formula into quadratic form, and find the number that will be the 1 solution, for example (x - 2)^2, (x - 3)^2, etc. You got this one!

OpenStudy (anonymous):

Are you still confused on the 2nd step? I know it can be a little confusing..

OpenStudy (anonymous):

I understand everything except for how you got x^2−4x+4=0

OpenStudy (anonymous):

Ok, well in order for an equation to have 1 solution, it must be like (x-2)^2, (x-3)^2, etc. Expand (x-2)^2 and tell me what you get.

OpenStudy (anonymous):

x^2-4x+4

OpenStudy (anonymous):

So, you see where I got the equation now?

OpenStudy (anonymous):

Yeah, I got it. So is the final answer for what does K equal 6?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Thanks! Can you help me with one more problem? I'll post it now.

OpenStudy (anonymous):

Yeah sure!